# Communication-Efficient Federated Fine-Tuning of Language Models via Dynamic Update Schedules

### v1

@@ 7. CONCLUSION
Appendix A EXPERIMENTAL APPROACH
A.1. Datasets & Task Descriptions
In this section, we describe the datasets used in our experiments, each corresponding to a distinct NLU task.
MRPC. The Microsoft Research Paraphrase Corpus (MRPC) (Dolan and Brockett, 2005) consists of sentence pairs collected from news-wire articles. The goal is to classify whether or not a given pair of sentences are paraphrases of each other. The reported metric is Accuracy. MRPC belongs to the category of Sentence Similarity tasks. Notably, the dataset is imbalanced, with 68% of the pairs labeled as paraphrases.
SST-2. The Stanford Sentiment Treebank (SST-2) (Socher et al., 2013) consists of sentences from movie reviews along with human annotations of their sentiment. The goal is to predict the sentiment of a given sentence as either positive or negative. The reported metric is Accuracy. SST-2 belongs to the category of Sentiment Analysis tasks.
RTE. The Recognizing Textual Entailment (RTE) (Wang et al., 2018) consists of pairs of sentences where the task is to determine whether or not the meaning of the second sentence (the hypothesis) is entailed by the first sentence (the premise). The reported metric is Accuracy. RTE belongs to the category of Natural Language Inference (NLI) tasks.
QNLI. The Question-answering Natural Language Inference (QNLI) dataset (Wang et al., 2018) consists of question-context pairs. The goal is to determine whether or not the context sentence contains the answer to the question. The reported metric is Accuracy. QNLI belongs to the category of NLI tasks.
MNLI m/mm. The Multi-Genre Natural Language Inference (MNLI) corpus  (Williams et al., 2018) is a large-scale dataset for evaluating a model’s ability to perform NLI across diverse textual domains. Each sample consists of a premise and a hypothesis, and the goal is to classify their relationship as entailment, contradiction, or neutral. The evaluation is conducted on two test sets: matched (MNLI-m) and mismatched (MNLI-mm), representing in-domain and out-of-domain genres, respectively. The reported metric is Accuracy.
A.2. Label Imbalance Scheme
Modeling label imbalance in a way that closely mirrors real-world scenarios has been extensively studied in the literature. The widely adopted approach is to simulate non-IID label distributions using a Dirichlet allocation scheme. The method works as follows (Li et al., 2022; Lin et al., 2022):
Assume we have a dataset with $L$ labels. The label distribution for each client is parameterized by a vector ${\bm{q}}^{(k)}\in(0,1)^{L}$ with $||{\bm{q}}^{(k)}||_{1}=1$, where ${\bm{q}}^{(k)}_{l}$ represents the proportion of instances of label $l$ for client $k$. We sample ${\bm{q}}^{(k)}$ from a Dirichlet distribution, ${\bm{q}}^{(k)}\sim Dir_{L}(\alpha{\bm{p}})$, where ${\bm{p}}$ is the prior class distribution (assumed to be known) and $\alpha>0$ is a concentration parameter. After constructing these sampled label distributions, we allocate data to each client accordingly.
The degree of label imbalance can be flexibly controlled by varying a parameter $\alpha$. Larger values of $\alpha$ result in more uniform distributions, with $\alpha\to\infty$ producing a perfectly uniform distribution across clients. Conversely, as $\alpha$ approaches zero, the distributions become highly imbalanced, with each client predominantly holding samples from a single label.
In our experiments we use a concentration parameter $\alpha=1.0$. Figure 6 visualizes its effect on label imbalance across clients for all datasets. Moreover, this figure also showcases the slight quantity imbalance which is inevitably introduced following aforementioned scheme.
Figure 6. Histogram of Label and Quantity data distribution across clients for all datasets. The non-IID label partitioning is done using a Dirichlet distribution with parameter $\alpha=1.0$
A.3. Hyper-parameter Selection
@@
In this section we outline the approach for identifying the optimal configuration for each FedOpt algorithm when training a specific model on a specific dataset. Algorithms in the FedOpt family are highly sensitive to both client and server learning rates, and their interaction is often task-specific and unpredictable. We conduct a grid search to determine the best hyper-parameters for each dataset and algorithm. The final selections based on this process are summarized in Table 5.
@@
As a representative example, consider the RoBERTa model trained MRPC dataset. We tune learning rates using a commonly adopted search grid (Wang et al., 2021) (Figure 7). Specifically we investigate server and client learning rates in $\{10^{-5},10^{-4},\dots,10^{-1}\}$ for FedAdaGrad, following well-established guidelines (Duchi et al., 2011). We then train RoBERTa with FedAdaGrad for $T=100$ rounds, recording the highest accuracy achieved across all rounds. As shown in Figure 7, the best-performing configuration for FedAdaGrad on MRPC reached $89\%$ accuracy with server and client learning rates of $10^{-4}$ and $10^{-5}$, respectively. These values are taken as the best hyper-parameters to train RoBERTa with FedAdaGrad on MRPC.
@@
Figure 7. The highest testing accuracy of RoBERTa on the MRPC dataset during the first 100 rounds of training, across various client and server optimizer learning rates. The MRPC test set is imbalanced, with $68.4\%$ of the pairs labeled as positive and $31.6\%$ as negative. As a result, these two accuracy values indicate that the model is always choosing one class, and should be treated as signs of non-convergence. Notably, FedAvg has an averaging aggregation scheme.
Table 5. The best-performing learning rate (lr) in $\log_{10}$ for the RoBERTa model with each FedOpt algorithm across datasets. ClientOpt is listed on top and ServerOpt at the bottom. For example, the bottom-left cell shows AdaGrad at the server with lr $=10^{-4}$ and SGD at the client with lr $=10^{-5}$
MRPC RTE SST-2 QNLI MNLI
FedAvg SGD 00 00 00 00 00
SGD $-3$ $-3$ $-3$ $-3$ $-2$
FedAvgM SGDM 00 $-1$ 00 00 00
SGD $-3$ $-3$ $-4$ $-3$ $-3$
FedAdam Adam $-4$ $-5$ $-4$ $-4$ $-4$
SGD $-4$ $-5$ $-4$ $-3$ $-2$
FedAdamW AdamW $-4$ $-5$ $-4$ $-4$ $-4$
SGD $-5$ $-5$ $-4$ $-3$ $-4$
FedAdaGrad AdaGrad $-4$ $-4$ $-4$ $-4$ $-4$
SGD $-5$ $-3$ $-4$ $-3$ $-2$

### v2

@@ VII Conclusion
-A Hyper-parameter Selection
@@
In this Appendix, we detail the approach for identifying the optimal configuration for each FedOpt algorithm when training a specific model on a specific dataset. Algorithms in the FedOpt family are highly sensitive to both client and server learning rates, and their interaction is often task-specific and unpredictable. We conduct a grid search to determine the best hyper-parameters for each dataset and algorithm. The final selections based on this process are summarized in Table V.
@@
As a representative example, consider the RoBERTa model trained MRPC dataset. We tune learning rates using a commonly adopted search grid [33] (Figure 6). Specifically we investigate server and client learning rates in $\{10^{-5},10^{-4},\dots,10^{-1}\}$ for FedAdaGrad, following well-established guidelines [18]. We then train RoBERTa with FedAdaGrad for $T=100$ rounds, recording the highest accuracy achieved across all rounds. As shown in Figure 6, the best-performing configuration for FedAdaGrad on MRPC reached $89\%$ accuracy with server and client learning rates of $10^{-4}$ and $10^{-5}$, respectively. These values are taken as optimal for training RoBERTa with FedAdaGrad on MRPC.
@@
Figure 6: The highest testing accuracy of RoBERTa on the MRPC dataset during the first 100 rounds of training, across various client and server optimizer learning rates. The MRPC test set is imbalanced, with $68.4\%$ of the pairs labeled as positive and $31.6\%$ as negative. As a result, these two accuracy values indicate that the model is always choosing one class, and should be treated as signs of non-convergence. Notably, FedAvg has an averaging aggregation scheme
TABLE V: Best-performing learning rates ($\log_{10}$) for training RoBERTa and DeBERTaV3 on different datasets using each FedOpt algorithm. ServerOpt is shown on top and ClientOpt on the bottom within each cell. For example, the bottom-left cell corresponds to training RoBERTa on MRPC with FedAdaGrad, where the optimal learning rates are $10^{-4}$ for server-side AdaGrad and $10^{-5}$ for client-side SGD
RoBERTa DeBERTaV3
MRPC RTE SST-2 QNLI MNLI-m/mm
FedAvg 00 00 00 00 00
$-3$ $-3$ $-3$ $-3$ $-2$
FedAvgM 00 $-1$ 00 00 00
$-3$ $-3$ $-4$ $-3$ $-3$
FedAdam $-4$ $-5$ $-4$ $-4$ $-4$
$-4$ $-5$ $-4$ $-3$ $-2$
FedAdamW $-4$ $-5$ $-4$ $-4$ $-4$
$-5$ $-5$ $-4$ $-3$ $-4$
FedAdaGrad $-4$ $-4$ $-4$ $-4$ $-4$
$-5$ $-3$ $-4$ $-3$ $-2$